# LEAM: A Prompt-only Large Language Model-enabled Antenna Modeling Method

### v1

@@ IV Conclusion
LEAM, an open-source, prompt-only LLM-enabled antenna modeling technique designed for automated antenna modeling, has been presented. Using three cases with complete description, weak description, and multi-modal description (i.e., image and text scanned for the literature), the effectiveness of LEAM has been shown. Particularly, LEAM realized prompt-only (fine-tuning and retrieval-augmented generation are not needed). The effectiveness comes from the proposed systematic framework and the defined tools. To the best of our knowledge, LEAM is the first LLM-based antenna modeling tool. Future works include enhancing LEAM’s compatibility with more LLMs (e.g., DeepSeek, Llama) and EM simulation software (e.g., Ansys HFSS, MATLAB Antenna Toolbox).

### v2

@@ IV Conclusion
In this paper, we present an LLM-based intelligent antenna design system, including an open-source LLM-enabled antenna modeling (LEAM) tool and an LLM-controlled optimization tool. Using an example of modeling and optimizing an antenna from the literature, the effectiveness of LADS has been shown. The effectiveness comes from the proposed systematic framework and the defined tools, including the LLM tools and the optimization tool SB-SADEA in [liu2022efficient]. To the best of our knowledge, LADS is the first attempt to jointly apply LLMs for antenna modeling and optimization. Future works include improving LADS’s compatibility with more LLMs (e.g., DeepSeek, Llama) and EM simulation software (e.g., Ansys HFSS, MATLAB Antenna Toolbox). In addition, fine-tuning LLMs on domain-specific data could further enhance the LADS’s usability in complex antenna design scenarios.